Two non negative integers are chosen at random.The probabability that their sum is divisible by 5 is
a.7/25
b.8/25
c.9/25
d.2/5

Grade:12

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
GivenTwo non negative integers are chosen at random.
Let the 2 non -ve integers be a,b then 0<=a,b<=9
Number of ways =10*10=100
The probabability that their sum is divisible by 5 is=
There are 20favorable cases like (0,0)(0,5)(1,4)(2,3)(3,2)(4,1)(5,0)
(1,9)(2,8)(3,7)(4,6)(5,5)(6,4)(7,3)(8,2)(9,1)
(7,8)(8,7)(9,6)(6,9)
So required probability=20/100=1/5